130% of what number is 169?
Found 2 solutions by flame8855, ojasvisunita:
Answer by flame8855(424)

130x/100= 169
x = 130

Answer by ojasvisunita(5)

Suppose the required number is A
130% of A=169
130/100*A=169
A=169*100/130
A=130